OpenStudy (anonymous):

What is the distance between the points (8, 3) and (8, –6)? A. 3 units B. 8 units C. 9 units D. 17 units

jimthompson5910 (jim_thompson5910):

Are you able to plot the two points?

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm d=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2} }\)

OpenStudy (e.mccormick):

Well, you can use the dystance formula, but there is something you should notice about those points that makes it even easier.

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 8}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 8}}\quad ,&{\color{blue}{ -6}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

I don't get it

jimthompson5910 (jim_thompson5910):

Here's what you would get if you plotted the two points on a xy grid

jimthompson5910 (jim_thompson5910):

What do you notice about the points?

OpenStudy (e.mccormick):



OpenStudy (anonymous):

i get it now

jimthompson5910 (jim_thompson5910):

so what is the distance between the two points?

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (e.mccormick):

WooWoo! =) Made it through the lag and did good.

OpenStudy (anonymous):

Thank you for your guys help i appreciate it!